OpenStudy (anonymous):

Find the value. 3/64 A. 2 B. 4 C. 8 D. 262,144

OpenStudy (anonymous):

@emmathis

OpenStudy (anonymous):

wiil give medal and fan

OpenStudy (anonymous):

Divide?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

3/64=3/64 is cannotbe simplified.

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (anonymous):

@emmathis

OpenStudy (anonymous):

pls help me

OpenStudy (studygurl14):

Okay, so \(\large\sqrt[3]{64}\)

OpenStudy (anonymous):

yes

OpenStudy (studygurl14):

This is the same thing as writing \(\large 64^{\frac{1}{3}}\)

OpenStudy (studygurl14):

Basically, what you're trying to find is a value, that when multiplied against itself three times, equals 64. Does that make sense?

OpenStudy (anonymous):

no

OpenStudy (studygurl14):

Okay. So...for example. \(\large\sqrt[3]{8}=2\) Because... \(\large 2^3 = 8\) Now does it make sense?

OpenStudy (anonymous):

so the answer is 8

OpenStudy (anonymous):

????????????

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (studygurl14):

No...I didn't say that.

OpenStudy (studygurl14):

I was just giving an example

OpenStudy (studygurl14):

What number, when multiplied by itself three times, equals 64?

OpenStudy (anonymous):

i dont know

OpenStudy (studygurl14):

Try. Just plug in some numbers. \(\large 3^3 = (3)(3)(3) = 9(3) = 27\) \(\large 27\neq 64\), so 3 is NOT the answer. Since 27 is smaller than 64, you know that the answer is \(\large\sqrt[3]{64}\) has to be bigger than 3.

OpenStudy (anonymous):

A. 2 B. 4 C. 8 D. 262,144

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (studygurl14):

Well...since 3 doesn't work, neither will 2, because it is too small. We already ruled out 8. and D is way too big. So...what answer choice is left?

OpenStudy (anonymous):

oh i get it know i was being stupid:) THANKS

OpenStudy (studygurl14):

You're welcome. :)